OpenStudy (yinshy):

A circle has a radius of 5 cm. What is the approximate area of the circle? A. 157.0 cm2 B. 19.6 cm2 C. 39.3 cm2 D. 78.5 cm2

OpenStudy (kendricklamar2014):

\[A = \pi \times r^2\]

OpenStudy (kendricklamar2014):

Now, just substitute :)

OpenStudy (yinshy):

so b?

OpenStudy (kendricklamar2014):

Nope

OpenStudy (yinshy):

C

OpenStudy (curiousquestions):

A=(3.14)*5^2 A=(3.14)*(5*5)

OpenStudy (kendricklamar2014):

No, what is 5^2

OpenStudy (curiousquestions):

5 to the second

OpenStudy (yinshy):

10

OpenStudy (kendricklamar2014):

\[5^2 = 5 \times 5 = 25\]

OpenStudy (curiousquestions):

5^2=5*5 What is 5 * 5?

OpenStudy (kendricklamar2014):

Do you understand that? @Yinshy

OpenStudy (curiousquestions):

A=3.14 * 25

OpenStudy (yinshy):

lol srry always think it means 5X2

OpenStudy (yinshy):

D

OpenStudy (kendricklamar2014):

Yes, thats correct

OpenStudy (curiousquestions):

Correct

OpenStudy (yinshy):

medal? me?

OpenStudy (kendricklamar2014):

I gave a medal to both of you :D

OpenStudy (yinshy):

oh ty

OpenStudy (kendricklamar2014):

My pleasure

OpenStudy (curiousquestions):

Thx for the medal! :)